OpenStudy (drakedoesmc):

A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking you select a sock at random replace it and select a second sock at random what s the probability that the first sock is blue and the second sock is red A. 3/5 B. 7/20 C. 3/25 D. 7/100

OpenStudy (mathstudent55):

Since these are two independent events, find the probability of each event, and then multiply the probabilities together.

OpenStudy (drakedoesmc):

so it would be 3/10 * 4/10= 7/100

OpenStudy (drakedoesmc):

right.

OpenStudy (mathstudent55):

Be careful with the multiplication of fractions. (Multiply the numerators and multiply the denominators.) \(\dfrac{a}{b} \times \dfrac{c}{d} = \dfrac{ac}{bd} \) \( \dfrac{3}{10} \times \dfrac{4}{10} = \dfrac{12}{100} = \dfrac{3}{25} \) (You added the numerators and multiplied the denominators.)